OpenStudy (marissalovescats):

AP Calc BC help! I recently took my test and I'm doing my test corrections. I originally got this problem wrong and I was very surprised. If someone could check my new answer, I'd greatly appreciate it. (Warning it's a somewhat lengthy problem)

OpenStudy (marissalovescats):

Find f '' (x): \[f(x) = \sqrt{2x^2+5}\]

OpenStudy (marissalovescats):

Without simplifying it (not being sure if I can, also since I'm kinda lazy) I got a very messy answer of: \[f '' (x) = \frac{ 8\sqrt{2x^2+5}-16x^2 }{ 8x^2+20[\sqrt{2x^2+5]} }\]

OpenStudy (marissalovescats):

I know my first derivative is right because that's just a basic chain rule. But the 2nd is maybe iffy. f '(x)= 2x/(2x^2+5)^1/2

OpenStudy (phi):

did you use the quotient rule or product rule for the second derv.

OpenStudy (marissalovescats):

I used chain and quotient

geerky42 (geerky42):

To make second derivative easier, you can just do this: \(f'(x) = \dfrac{2x}{\sqrt{2x^2+5}}\Rightarrow f'(x) = 2x(2x^2+5)^{-\frac{1}{2}}\) Use product rule from here

OpenStudy (marissalovescats):

Yeah I realize that but idk why I didn't do it that way, I just perfer the quotient in a situation like this

OpenStudy (marissalovescats):

I'll do it like that one sec

OpenStudy (marissalovescats):

I got something I don't wanna simplify lol

OpenStudy (phi):

For the top I get \[ 2 (2x^2+5)^{\frac{1}{2}} - x^2 (2x^2+5)^{-\frac{1}{2}} \]

OpenStudy (phi):

and the bottom is 2x^2+5

OpenStudy (marissalovescats):

With the chain and quotient I got [4(2u^1/2) - 4x (u^1/2)] 4x all over (2u^1/2)^2

OpenStudy (marissalovescats):

Since I didn't simplify f ' a first and kept 4x/2(2x^2+5)^1/2

OpenStudy (marissalovescats):

the 2nd one would be ^1/2 in there

OpenStudy (marissalovescats):

-1/2 jeez I cant even breathe right today

geerky42 (geerky42):

Using quotient rule, I got \(f''(x)=\dfrac{2\sqrt{2x^2+5}-4x^2(2x^2+5)^{-\frac{1}{2}}}{2x^2+5}\)

OpenStudy (phi):

d u/v = (v du - u dv)/v^2 in this case u = 2x du = 2 v= sqr(2x^2+5) dv = 2x/sqr(2x^2+5)

OpenStudy (phi):

I don't think you are doing the rule correctly, but I can't tell what you did.

OpenStudy (marissalovescats):

I am writing it all out now and going to post a picture so you guys can see how I did it.

OpenStudy (phi):

\[ \frac{2 (2x^2+5)^{\frac{1}{2}} (2x^2+5)^{\frac{1}{2}} - 4x^2 (2x^2+5)^{-\frac{1}{2}}(2x^2+5)^{\frac{1}{2}} }{(2x^2+5)^\frac{3}{2}} \\ =\frac{2(2x^2+5) - 4x^2}{(2x^2+5)^\frac{3}{2}} \\ = \frac{10}{(2x^2+5)^\frac{3}{2}} \]

OpenStudy (marissalovescats):

used the chain rule and then the product rule.

OpenStudy (marissalovescats):

Agreed?

OpenStudy (phi):

yes

OpenStudy (marissalovescats):

Then I plug u back in to get: \[\frac{ 2 }{ (2x^2+5)^{1/2}} - \frac{ 4x^2 }{ (2x^2+5)^{3/2} }\] Agreed?

OpenStudy (marissalovescats):

Then I get a common denominator by doing (2x^2+5)^1/2 times (2x^2+5)^3/2 Which gives me (2x^2+5)^2 as my denominator.

OpenStudy (phi):

yes, and if you multiply the first term by u/u it simplifies

OpenStudy (phi):

your common denominator will be \( (2x^2+5)^\frac{3}{2} \)

OpenStudy (phi):

or least, that will be the LCD

OpenStudy (marissalovescats):

No it's (2x^2+5)^2 because you multiply the 2 denominators together for a common and 1/2 +3/2 = 4/2 which is 2

OpenStudy (phi):

that is messy. use (stuff)^(3/2)

OpenStudy (marissalovescats):

And I muliply the left by (2x^2+5)^3/2 and the right by (2x^2+5)^1/2 To get a final answer of: 2(2x^2+5)^3/2 - 4x^2(2x^2+5)^1/2 all over (2x^2+5)^2

OpenStudy (marissalovescats):

Simplifying is my enemy, honestly.

OpenStudy (phi):

you could and it's not wrong. but notice \[ \frac{ 2(2x^2+5) }{ (2x^2+5)^{1/2}(2x^2+5) } - \frac{ 4x^2 }{ (2x^2+5)^{3/2} } \]

OpenStudy (marissalovescats):

You have an extra (2x^2+5)

OpenStudy (phi):

I multiplied by top and bottom

OpenStudy (marissalovescats):

?

OpenStudy (marissalovescats):

Oh I see what you're doing.

OpenStudy (marissalovescats):

You're doing that so you'll just have (2x^2+5)^3/2

OpenStudy (phi):

This was your result \[ \frac{ 2 }{ (2x^2+5)^{1/2}} - \frac{ 4x^2 }{ (2x^2+5)^{3/2} } \] multiply the first term now you can simplify the top

OpenStudy (marissalovescats):

10/(2x^2+5)^2/3

OpenStudy (phi):

^3/2 but yes, that looks a little better.

OpenStudy (marissalovescats):

I see what you did, yeah that's a lot easier. I just suck at simplifying lol

OpenStudy (phi):

It is like a lot of things. Practice.

OpenStudy (marissalovescats):

Yeah definitely agree lol

OpenStudy (marissalovescats):

Thanks so much for helping, I have a couple others in a minute that are different so if you see me post maybe you can help again lol